tutors I really need you all to explain to me how to solve this problem
which of the following are factors of P(x)= x^3 - 5x^2 + 3x + 9
a. x-1
b. x+3
c. x-3
I would appreciate any help I can get!

Answer by edjones(8007)

P(x)= x^3 - 5x^2 + 3x + 9
.
If x-1 is a factor then x=1 and 1-5+3+9=0 false.
If x+3 then x=-3 and -27-45-9+9=0 false.
If x-3 then x=3 and 27-45+9+9=0 true.
.
Ed